OpenStudy (anonymous):

Solve for M A=2(m-w)

OpenStudy (anonymous):

m= a+w??? Is this correct?

OpenStudy (anonymous):

@Jboy729 a = 2 (m - w) 2m - 2w = a 2m - 2w + 2w = a + 2w 2m = a + 2w 2m/2 = a + 2w/2 m = 1/2a + w

OpenStudy (anonymous):

m=a/2 +w???

OpenStudy (anonymous):

Thanks!